# Finite Dynamical Laminations

### v2

@@ Abstract
We developed several combinatorial notions about laminations some with clear implications for parameter space. We introduce a simplified class of laminations called finite dynamical laminations (FDL). In order to count FDL, we introduce sibling portraits, of which we provide a comprehensive counting theorem. We provide a characterization of which periodic polygons appear in invariant laminations. We introduce the pullback tree. The base of the pullback tree is a set of laminations, and we show that those laminations are proper and invariant and have a restriction on its critical leaves, so all laminations in the base of the pullback tree correspond to a polynomial. We define the generational FDL graph, and it provides a summary of the combinatorial information we provide about polynomial parameter space.
This work was supported by a UAB Honors College Presidential Honors Fellowship. The author’s research advisor was Dr. John C. Mayer. Some revisions were made under the advice of Dr. Lex G. Oversteegen.
@@ 1 Introduction
A lamination is a closed set of chords of the unit disk that do not cross each other except at the endpoints, and contains all chords that consist of a single point on the boundary of the circle. A leaf is a chord in a lamination joining 2 different points while a degenerate leaf is a point of the circle.
@@
The dynamical ray is the image of the straight ray under $\Psi$. The dynamic ray lands at a point in the Julia set. The lamination of a polynomial has the equivalence relation such that two angles are in the same class iff their dynamical rays land at the same point. This explanation is illustrated in a 5-minute video [2]. Dropping the assumption that the Julia set is connected and locally connected, the definition of dynamical ray can be based on a potential function around the filled Julia set [3], but it is no longer the case that all rays land, and it is harder to associate laminations with polynomials.
Laminations have been used to develop an exceedingly powerful and detailed model of the Mandelbrot set. However, it is much harder in higher degree. In the case of cubics, there is uncertainty about the shape of the set of polynomials associated with the empty lamination. Indeed, the sheer dimension of the set of polynomials makes it difficult to understand the relationships between the polynomials of invariant laminations. Therefore, we introduce a set of simple laminations with a straightforward connection to parameter space, and we aim for a less detailed description of parameter space.
This document explores the uses of finite laminations for studying polynomials. Section 2 introduces the terminology for discussing such laminations, and proves which finite laminations are part of an invariant lamination. In Section 3, we count sibling portraits, which we introduced in Section 2. In Section 4, we introduce Finite Dynamical Lamination, or FDL, a definition which helps us extend combinatorial questions and helps us study the structure of parameter space. Section 5 discusses the limits of sequences of FDL, and Section 6 proves which that all FDL are realized by a polynomial. Section 7 finally shows how to assemble graphs that represent parts of parameter space.
@@ 2 The Critical Chords Problem
Given a finite set of disjoint, periodic polygons, when does there exist a $\sigma_{d}$-invariant lamination that contains them? From [4], it is known that the question is the same as whether there is a set of $d-1$ disjoint critical leaves for that set of polygons. Our question is answered in [5], but the argument is longer than necessary. Unlike that document, we will find the critical chords directly, and leave it to the reader to consider how to create an invariant lamination using those leaves.
@@
A $\sigma_{d}$-invariant lamination should be thought of as a $d$ to $1$ invariant lamination. The reader may use either the definition of invariant lamination found in [1] or in [4]. Less than invariant, a lamination may be called forward invariant meaning that the lamination contains its $\sigma_{d}$-image. The next theorem is well known. In addition to proposition II.4.5 of [4], Dr. Oversteegen has promised to write up a proof.
@@
Given $d-1$ critical leaves compatible with a forward invariant lamination, there exists an invariant lamination containing the original that is also compatible with the critical chords.
@@
A round gap is the closure of a component of the complement of a lamination that has arcs of the circle in its boundary. The empty lamination contains one round gap.
@@
Since we mostly think about polygons of a lamination, as opposed to its leaves, and those polygons have siblings, we give a name to a set of sibling polygons. Since a round gap with a degree is like the whole disk, this definition can be with reference to a round gap or the whole disk. We define sibling portrait in a way that is not consistent with the existing notion of sibling collection: a sibling portrait of a polygon $P$ is not, in general, composed of $P$’s siblings.
@@
Since invariant laminations have siblings for every polygon, the question of whether we can create an invariant lamination comes down to whether we can find places for the siblings. That is, each polygon must be in some sibling portrait that could be added to the lamination. We work to show when this is possible. The next few proofs rely on the order of vertices in a polygon. This is the counterclockwise order of the vertices. When indexing the vertices of an n-gon, subscript is implied to be modulo n.
@@
Since polygons are disjoint even at their endpoints, each preimage vertex is, at most, part of two sides of one polygon. If the original polygon has an order of vertices like $x_{1}x_{2}x_{3}\dots x_{n}$ then each $x_{k}$ must be used the maximal number of times in preimages of $\overline{x_{k-1}x_{k}}$ and $\overline{x_{k}x_{k+1}}$. If we have a 2-gon, then we have an order $ab$ and in that case, each vertex must be used in two sides of a polygon which map to the two sides of the original polygon. Consider a walk around the sides of a polygon. We cannot afford to visit vertices in an order other than that of the original polygon. This excludes not only leaves whose image are interior leaves, but also critical leaves. Moreover, we can only return to our starting place if we are at the end of a cycle. ∎
@@
The preimages of the vertices of a polygon are evenly spaced throughout the circle in the same repeating order as the polygon. (To see this, consider pulling back the polygon according to an all critical d-gon; a half open arc under any side of the d-gon maps monotone onto the circle.) For this reason, there is always at least on sibling portrait that can be formed by using all the vertices in one polygon. If a polygon maps with a degree (with positive orientation) then its siblings are in the same repeating order as itself.
@@
Label the vertices of the original polygon, $\sigma_{d}(P)$, as $x_{1}x_{2}x_{3}\dots x_{n}$. Since the $\sigma_{d}(P)$ has a leaf from each $x_{k}$ to each $x_{k+1}$; $P$ must also. Therefore, as we walk around the leaves of $P$, we find vertices in a, possibly repeating, order of either $x_{1}x_{2}x_{3}\dots x_{n}$ or $x_{n}x_{n-1}x_{n-2}\dots x_{1}$. Since a polygon is a convex hull, the order of its vertices that we get by walking its leaves is the same as the order that they appear on the circle.
Suppose that we start drawing $P$ one leaf at a time starting with $x_{1}$, and draw the leaves in CCW, counterclockwise, order. In all of $R_{j}$, we have the preimages of the vertices in order like $x_{1}x_{2}x_{3}\dots x_{n}x_{1}x_{2}x_{3}\dots x_{n}\dots$. Suppose that we then connect from $x_{k}$ to $x_{k-1}$ while moving CCW by making $P$ orientation reversing. Then the open arc under that leaf would contain $x_{k+1}$ and $x_{k-2}$ the start and end respectively. Using the repeating order of vertices in the arc, we see that it contains fewer $x_{k}$ vertices than $x_{k+1}$. Thus, we can not form complete groups of vertices as needed in the previous lemma.
@@
The following lemma is the contribution of Dr Oversteegen. The proof does not require that $\overline{xy}$ is non-crossing with $\overline{ab}$, but it does require that they not be siblings.
@@
Consider that the siblings of $a$,$b$,$x$,$y$ are evenly spaced throughout the circle in a fixed order. We label the siblings of these 4 points without distinguishing one sibling from another. If the $x$ or $y$ are equal to some sibling of $a$ or $b$, then their images obviously cannot be disjoint. Assume without loss of generality that $x$ is in some arc bounded by $a$ and then $b$. If $y$ is in an interval bounded by a sibling of first $b$ and then $a$, then the repeating order of the 4 points is $axby\dots$. A half open arc starting and ending with $a$ will map monotone onto the circle. Thus, the image of $\sigma_{d}(\overline{xy})$ crosses $\sigma_{d}(\overline{ab})$. ∎
@@
As long as there is a polygon, $P$, left that is not part of a sibling portrait, we can add a polygon that is its sibling. Consider a sibling $a^{\prime}$ of a vertex of $P$. $a^{\prime}$ is on the boundary of a round gap, call it $R$. Every sibling of a vertex $P$ is either in an open arc of $R$ or in the closure of one of its holes. These holes fall into one of 2 categories: either they are bounded by a leaf that is part of a polygon sibling to $P$, or it isn’t. In the first case Lemma 2.10 applies to that polygon; thus, on the side of the polygon containing $R$, there is a complete set of vertices. Using Lemma 2.11, we can make the same conclusion about the other type of hole. Each closure of a hole contains an equal number of siblings of each vertex/letter. Since each vertex has the same number of siblings in the whole disk, each vertex has the same number of siblings in the union of the open arcs of $R$.
@@
    If $R_{j}$ is a round gap with of degree $d_{j}$, and $P$ is a polygon contained in $\sigma_{d}(R_{j})$ then the round gaps resulting from inserting a sibling portrait into $R_{j}$ each have a degree.
@@
    Consider the set of gaps (including polygons) $\{G_{i}\}$ formed when inserting that sibling portrait into $R_{j}$, and their degrees ${d_{i}}$. Then
    $\sum_{i}(d_{i}-1)=d_{j}-1$
@@
    Moreover, given a lamination that consists of finitely many disjoint polygons such that each polygon is part of a sibling portrait in the whole disk, all gaps have a degree and the equation above holds. In this case $d=d_{j}$.
@@
    By Lemma 2.10 after some “first” polygon is inserted, a critical arc created must have, up to starting point, have the same repeating order as $R_{j}$. Let the start of the arc be $x_{k}$. Thus, its end is $x_{k+1}$. Moreover, the first preimage of a vertex appearing in the open arc is also $x_{k+1}$. We can also note that the sub-arc at the beginning and ending of the critical arc are siblings as well as being consecutive in the circular order of the arc. We then add a polygon starting with $x_{k+1}$ since that is the first preimage vertex of the open arc, and thus the last point of the polygon is $x_{k}$ by the polygon’s circular order. In this way, as we remove from the arc by adding polygons until there are no critical arcs, what remains is a number (at least 2) of sibling arcs of the circle. These arcs and leaves of the first and subsequent polygons form the boundary of a round gap. From the parent polygon, we take the leaf $\overline{x_{k}x_{k+1}}$ and from every child polygon, we also take the leaf $\overline{x_{k}x_{k+1}}$. Since the arcs in-between the leaves of this gap all have the same end points, they are also siblings. Therefore, the round gap that is formed has some degree equal to the number of polygons in its boundary.
@@
    Consider working up to the lamination inductively, but one sibling portrait at a time. The statement is true of the empty lamination. At each step while inserting a sibling portrait, in-between other sibling portraits, we can see it as inserting a sibling portrait into a round gap with say degree $d_{j}$. Since a round gap with a degree is like the disk, the equation holds for the new gaps inside the round gap. Since the new terms of the sum are equal to what they replace, the equation holds.
The formula above implies something comparable to the Riemann-Hurwitz relation found in [6] when it is applied to the polygons and the round gaps of the lamination. That is the criticality trapped in the polygons plus the criticality free in the round gaps adds to $d-1$.
After forming the sibling portraits, we can place $d-1$ critical chords (without forming a polygon out of them) into gaps in at least one way. Any point in the basis of a critical gap could be the starting point of such a concatenation of critical chords. These critical chords separate the circle into $d$ branches of the inverse which by Theorem 2.3 creates an invariant lamination containing the original set of polygons.
@@
Another question is whether there exists an invariant lamination that contains every periodic orbit in some list and no more periodic polygons. This is obviously not true. For one example, in degree 3, take $L$ to be the rotational triangles of _001 and _112. The leaf from _0 to _1 is forced. In the original lamination, the round gap where the leaf is forced is not critical. For an example in degree 2, take $L$ to be the orbit of the leaf from _000101 to _101000. A rotational triangle is forced in the non-critical part of the critical gap. Perhaps there is some meaningful sense in which nothing is forced in a non-critical part of the original set.
@@ 3 Counting Sibling Portraits
Given a round gap $R$ of degree $i$ and an $n$-gon, $P$, contained in $\sigma_{d}(R)$, let $F(i,n)$ denote the number of sibling portraits, and let $f(i,n)$ denote the number of sibling portraits with all polygons one-to-one. We observe a bijection with the full rooted $n$-ary trees, which are counted in [7]. Though it is a function in two variables, that does not stop us from providing OEIS, [8], references: $f(i,2)$ = A000108$(i)$, $f(i,3)$ = A001764$(i)$, $f(i,4)$ = A002293$(i)$,$f(i,5)$ = A002294$(i)$. These are the Fuss–Catalan numbers.
@@
One-to-one sibling portraits of an $n$ sided polygon with in a degree $i$ round gap are in bijection with full, rooted, $n$-ary trees with $i$ internal vertices. Polygons correspond to internal vertices, and the trees are made according to the principle that the first CCW new polygon divides the region that it is in into $n$ smaller regions. In the bijection pictured in Figure 1, a depth first traversal of the internal vertices in the tree would correspond to a CCW ordering of the polygons based on their first ray. Such trees are known to be counted by the closed-form formula above [7].
@@
Let $B$ be the following function from the one-to-one portraits for $i$ and $n+1$ to the complete set of sibling portraits for $i$ and $n$. Take a sibling portrait from the domain. It is essentially a way of connecting $n+1$ colors each with $i$ points into disjoint polygons. Selecting a color arbitrarily, $X$. Collect the polygons into groups by transitively adding the polygon which appears after the $X$ vertex of any polygon. The output sibling portrait is a set of convex hulls of the sets of vertices other than $X$ from all the polygons in each of the groups.
@@
$B$ has the ascribed range because at neither step could it cause polygons to cross. It first merges polygons that are adjacent. Next, it removes vertices, which is replacing a polygon with a smaller polygon (the convex hull of a subset of the vertices) therefore, it will still be a lamination.
Without loss of generality, assume that the domain is the sibling portraits of a polygon with vertices in the positional order $abcd$. Let the range be sibling portraits of polygons of $abc$, and we correspond the vertices of each in the way suggested by the labeling, and of course $X=d$.
We form the inverse of $B$. By Lemmas 2.8 and 2.10, all the polygons of a sibling portrait map as a positively oriented covering map. Consider a polygon from an element of the range. It is in the form $abcabcabc\dots$ in positional order. In the whole circle there are multiple $d$ points that we could add, but we are forced to add the ones appearing right before the $a$ vertices of this polygon. (Expanding the polygon in that way cannot cause things to cross since it is the $d$ directly before the $a$). Next we need to break this into polygons, and the only way to do this consecutively ending with $d$. Suppose by way of contradiction that there is some $a$ vertex that we are allowed to connect a $d$ vertex to other than the $a$ preceding it, and finish that polygon as $P_{1}$. Consider an arc of the circle partitioned by that polygon and the vertices in that arc. At least one arc will have a sequence of vertices that does not start with $a$, and the set of polygons formed in that arc will not be placed with $P_{1}$ by $B$. Thus, there is exactly one sibling portrait of the domain for each of the range. ∎

### v3

@@ Abstract
We develop several combinatorial notions about laminations, some with clear implications for parameter space. We introduce a simplified class of laminations called finite dynamical laminations (FDL). In order to count FDL, we introduce sibling portraits, of which we provide a comprehensive counting theorem. We provide a characterization of which periodic polygons appear in invariant laminations. We introduce the pullback tree. The base of the pullback tree is a set of laminations, and we show that those laminations are proper and invariant and all laminations in the base of the pullback tree correspond to a polynomial. We define the generational FDL graph, and it provides combinatorial information about polynomial parameter space.
This work was supported by a UAB Honors College Presidential Honors Fellowship, and mostly written in 2024. The author’s research advisor was Dr. John C. Mayer. Some revisions were made under the advice of Dr. Lex G. Oversteegen.
@@ 1 Introduction
A lamination is a closed set of chords of the unit disk that do not cross each other except at the endpoints that additionally includes every point of the circle. A leaf is a chord in a lamination joining 2 different points, while a degenerate leaf is a point of the circle.
@@
The dynamical ray is the image of the straight ray under $\Psi$. The dynamic ray lands at a point in the Julia set. The lamination of a polynomial has the equivalence relation such that two angles are in the same class iff their dynamical rays land at the same point. This explanation is illustrated in a 5-minute video [2]. Dropping the assumption that the Julia set is connected and locally connected, a dynamical ray can be defined based on a potential function around the filled Julia set [3], but it is no longer the case that all rays land, and it is harder to associate laminations with polynomials.
Laminations have been used to develop an exceedingly powerful and detailed model of the Mandelbrot set. However, it is much harder in higher degrees. In the case of cubics, there is uncertainty about the shape of the set of polynomials associated with the empty lamination. Indeed, the sheer dimension of the set of polynomials makes it difficult to understand the relationships between the polynomials of invariant laminations. Therefore, we introduce a set of simple laminations with a straightforward connection to parameter space, and we aim for a less detailed description of parameter space.
This document explores the uses of finite laminations for studying polynomials. Section 2 introduces the terminology for discussing such laminations and proves which finite laminations are part of an invariant lamination. In Section 3, we count sibling portraits, which we introduced in Section 2. In Section 4, we introduce Finite Dynamical Lamination, or FDL, a definition that helps us extend combinatorial questions and helps us study the structure of parameter space. Section 5 discusses the limits of sequences of FDL, and Section 6 proves that all FDL are realized by a polynomial. Section 7 finally shows how to assemble graphs that represent parts of parameter space.
@@ 2 The Critical Chords Problem
Given a finite set of disjoint, periodic polygons, when does there exist a $\sigma_{d}$-invariant lamination that contains them? From [4], it is known that the question is the same as whether there is a set of d1d-1 disjoint critical leaves for that set of polygons. This section is dedicated to proving the existence of those critical cords when they exist.
@@
A lamination may be called forward invariant, meaning that the lamination contains its $\sigma_{d}$-image. The next theorem is well known. The following statement is proposition II.4.5 of [4].
@@
Given d1d-1 critical leaves compatible with a forward invariant lamination, there exists an invariant lamination containing the original that is also compatible with the critical chords.
@@
A round gap is the closure of a component of the complement of a lamination that has arcs of the circle in its boundary. The empty lamination has one round gap.
@@
Since we mostly think about polygons of a lamination, as opposed to its leaves, and those polygons have siblings, we give a name to a set of sibling polygons. Since a round gap with a degree is like the whole disk, this definition can be with reference to a round gap or the whole disk. We define sibling portrait in a way that is inconsistent with the existing notion of sibling collection: a sibling portrait of a polygon $P$ is typically not composed of $P$’s siblings.
@@
Since invariant laminations have siblings for every polygon, the question of whether we can create an invariant lamination comes down to whether we can find places for the siblings. That is, each polygon must be in some sibling portrait that could be added to the lamination. We work to show when this is possible. The next few proofs rely on the order of vertices in a polygon. This is the counterclockwise order of the vertices. When indexing the vertices of an $n$-gon, the subscript is implied to be modulo $n$.
@@
Since polygons are disjoint even at their endpoints, each preimage vertex is, at most, part of two sides of one polygon. If the original polygon has an order of vertices like $x_{1}x_{2}x_{3}\dots x_{n}$ then each $x_{k}$ must be used the maximal number of times in preimages of $\overline{x_{k-1}x_{k}}$ and $\overline{x_{k}x_{k+1}}$. If we have a 2-gon, then we have an order $ab$ and in that case, each vertex must be used in two sides of a polygon that map to the two sides of the original polygon. Consider a walk around the sides of a polygon. We cannot afford to visit vertices in an order other than that of the original polygon. This excludes not only leaves whose images are interior leaves but also critical leaves. Moreover, we can only return to our starting place if we are at the end of a cycle. ∎
@@
The preimages of the vertices of a polygon are evenly spaced throughout the circle in the same repeating order as the polygon. (To see this, consider pulling back the polygon according to an all-critical d-gon; a half-open arc under any side of the d-gon maps monotone onto the circle.) For this reason, there is always at least one sibling portrait that can be formed by using all the vertices in one polygon. If a polygon maps with a degree (with positive orientation), then its siblings are in the same repeating order as itself.
@@
Label the vertices of the original polygon, $\sigma_{d}(P)$, as $x_{1}x_{2}x_{3}\dots x_{n}$. Since the $\sigma_{d}(P)$ has a leaf from each $x_{k}$ to each $x_{k+1}$, $P$ must also. Therefore, as we walk around the leaves of $P$, we find vertices in a, possibly repeating, order of either $x_{1}x_{2}x_{3}\dots x_{n}$ or $x_{n}x_{n-1}x_{n-2}\dots x_{1}$. Since a polygon is a convex hull, the order of its vertices that we get by walking its leaves is the same as the order that they appear on the circle.
Suppose that we start drawing $P$ one leaf at a time starting with $x_{1}$ and draw the leaves in CCW, counterclockwise, order. In all of $R_{j}$, we have the preimages of the vertices in order like $x_{1}x_{2}x_{3}\dots x_{n}x_{1}x_{2}x_{3}\dots x_{n}\dots$. Suppose that we then connect from $x_{k}$ to $x_{k-1}$ while moving CCW by making $P$ orientation reversing. Then the open arc under that leaf would contain $x_{k+1}$ and $x_{k-2}$, the start and end, respectively. Using the repeating order of vertices in the arc, we see that it contains fewer $x_{k}$ vertices than $x_{k+1}$. Thus, we cannot form complete groups of vertices as needed in the previous lemma.
@@
The following lemma is the contribution of Dr. Oversteegen. The proof does not require that $\overline{xy}$ is non-crossing with $\overline{ab}$, but it does require that they not be siblings.
@@
Consider that the siblings of $a$, $b$, $x$, $y$ are evenly spaced throughout the circle in a fixed order. We label the siblings of these 4 points without distinguishing one sibling from another. If the $x$ or $y$ is equal to some sibling of $a$ or $b$, then their images obviously cannot be disjoint. Assume without loss of generality that $x$ is in some arc bounded by $a$ and then $b$. If $y$ is in an interval bounded by a sibling of first $b$ and then $a$, then the repeating order of the 4 points is $axby\dots$. A half open arc starting and ending with $a$ will map monotone onto the circle. Thus, the image of $\sigma_{d}(\overline{xy})$ crosses $\sigma_{d}(\overline{ab})$. ∎
@@
As long as there is a polygon, $P$, left that is not part of a sibling portrait, we can add a polygon that is its sibling. Consider a sibling $a^{\prime}$ of a vertex of $P$. $a^{\prime}$ is on the boundary of a round gap, call it $R$. Every sibling of a vertex $P$ is either in an open arc of $R$ or in the closure of one of its holes. These holes fall into one of 2 categories: either they are bounded by a leaf that is part of a polygon sibling to $P$, or they are not. In the first case Lemma 2.10 applies to that polygon; thus, on the side of the polygon containing $R$, there is a complete set of vertices. Using Lemma 2.11, we can make the same conclusion about the other type of hole. Each closure of a hole contains an equal number of siblings of each vertex/letter. Since each vertex has the same number of siblings in the whole disk, each vertex has the same number of siblings in the union of the open arcs of $R$.
@@
    If $R_{j}$ is a round gap with of degree $d_{j}$, and $P$ is a polygon contained in $\sigma_{d}(R_{j})$, then the round gaps resulting from inserting a sibling portrait into $R_{j}$ each have a degree.
@@
    Consider the set of gaps (including polygons) $\{G_{i}\}$ formed when inserting that sibling portrait into $R_{j}$, and their degrees $d_{i}$}. Then _i(d_i-1) = d_j-1.
@@
    Moreover, given a lamination that consists of finitely many disjoint polygons such that each polygon is part of a sibling portrait in the whole disk, all gaps have a degree, and the equation above holds. In this case $d=d_{j}$.
@@
    By Lemma 2.10 after some “first” polygon is inserted, a critical arc created must have, up to starting point, the same repeating order as $R_{j}$. Let the start of the arc be $x_{k}$. Thus, its end is $x_{k+1}$. Moreover, the first preimage of a vertex appearing in the open arc is also $x_{k+1}$. We can also note that the sub-arc at the beginning and ending of the critical arc are siblings as well as being consecutive in the circular order of the arc. We then add a polygon starting with $x_{k+1}$ since that is the first preimage vertex of the open arc, and thus the last point of the polygon is $x_{k}$ by the polygon’s circular order. In this way, as we remove from the arc by adding polygons until there are no critical arcs, what remains is a number (at least 2) of sibling arcs of the circle. These arcs and leaves of the first and subsequent polygons form the boundary of a round gap. From the parent polygon, we take the leaf $\overline{x_{k}x_{k+1}}$, and from every child polygon, we also take the leaf $\overline{x_{k}x_{k+1}}$. Since the arcs in between the leaves of this gap all have the same endpoints, they are also siblings. Therefore, the round gap that is formed has some degree equal to the number of polygons in its boundary.
@@
    Consider working up to the lamination inductively, but one sibling portrait at a time. The statement is true of the empty lamination. At each step while inserting a sibling portrait, in-between other sibling portraits, we can see it as inserting a sibling portrait into a round gap with, say, degree $d_{j}$. Since a round gap with a degree is like the disk, the equation holds for the new gaps inside the round gap. Since the new terms of the sum are equal to what they replace, the equation holds.
The formula above implies something comparable to the Riemann-Hurwitz relation found in [5] when it is applied to the polygons and the round gaps of the lamination. That is the criticality trapped in the polygons plus the criticality free in the round gaps adds to d1d-1.
After forming the sibling portraits, we can place d1d-1 critical chords (without forming a polygon out of them) into gaps in at least one way. Any point in the basis of a critical gap could be the starting point of such a concatenation of critical chords. These critical chords separate the circle into $d$ branches of the inverse, which by Theorem 2.3 creates an invariant lamination containing the original set of polygons.
@@
Another question is whether there exists an invariant lamination that contains every periodic orbit in some list and no more periodic polygons. This is sometimes not the case. For one example, in degree 3, take $L$ to be the rotational triangles of _001 and _112. The leaf from _0 to _1 is forced. In the original lamination, the round gap where the leaf is forced is not critical. For an example in degree 2, take $L$ to be the orbit of the leaf from _000101 to _101000. A rotational triangle is forced in the non-critical part of the critical gap.
@@ 3 Counting Sibling Portraits
Given a round gap $R$ of degree $i$ and an $n$-gon, $P$, contained in $\sigma_{d}(R)$, let $F(i,n)$ denote the number of sibling portraits, and let $f(i,n)$ denote the number of sibling portraits with all polygons one-to-one. We observe a bijection with the full rooted $n$-ary trees, which are counted in [6]. Though it is a function in two variables, that does not stop us from providing OEIS, [7], references: $f(i,2)$ = A000108(i)(i), $f(i,3)$ = A001764(i)(i), $f(i,4)$ = A002293(i)(i), $f(i,5)$ = A002294(i)(i). These are the Fuss–Catalan numbers.
@@
One-to-one sibling portraits of an $n$ sided polygon in a degree $i$ round gap are in bijection with full, rooted, $n$-ary trees with $i$ internal vertices. Polygons correspond to internal vertices, and the trees are made according to the principle that the first CCW new polygon divides the region that it is in into $n$ smaller regions. In the bijection pictured in Figure 1, a depth-first traversal of the internal vertices in the tree would correspond to a CCW ordering of the polygons based on their first ray. Such trees are known to be counted by the closed-form formula above [6].
@@
Let $B$ be the following function from the one-to-one portraits for $i$ and $n+1$ to the complete set of sibling portraits for $i$ and $n$. Take a sibling portrait from the domain. It is essentially a way of connecting $n+1$ colors, each with $i$ points into disjoint polygons. Selecting a color arbitrarily, $X$. Collect the polygons into groups by transitively adding the polygon that appears after the $X$ vertex of any polygon. The output sibling portrait is a set of convex hulls of the sets of vertices other than $X$ from all the polygons in each of the groups.
@@
$B$ has the ascribed range because at neither step could it cause polygons to cross. It first merges adjacent polygons. Next, it removes vertices, which is replacing a polygon with a smaller polygon (the convex hull of a subset of the vertices); therefore, it will still be a lamination.
Without loss of generality, assume that the domain is the sibling portraits of a polygon with vertices in the positional order abcdabcd. Let the range be sibling portraits of polygons of abcabc, and we correspond the vertices of each in the way suggested by the labeling, and of course $X=d$.
We form the inverse of $B$. By Lemmas 2.8 and 2.10, all the polygons of a sibling portrait map as a positively oriented covering map. Consider a polygon from an element of the range. It is in the form $abcabcabc\dots$ in positional order. In the whole circle there are multiple $d$ points that we could add, but we are forced to add the ones appearing right before the $a$ vertices of this polygon. (Expanding the polygon in that way cannot cause things to cross since it is the $d$ directly before the $a$). Next we need to break this into polygons, and the only way to do this consecutively is ending with $d$. Suppose by way of contradiction that there is some $a$ vertex that we are allowed to connect a $d$ vertex to other than the $a$ preceding it, and finish that polygon as $P_{1}$. Consider an arc of the circle partitioned by that polygon and the vertices in that arc. At least one arc will have a sequence of vertices that does not start with $a$, and the set of polygons formed in that arc will not be placed with $P_{1}$ by $B$. Thus, there is exactly one sibling portrait of the domain for each of the range. ∎